OpenStudy (rock_mit182):

To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationist's estimate of the size of the bear population?

OpenStudy (rock_mit182):

@mathmate @sourwing

OpenStudy (rock_mit182):

@Whitemonsterbunny17 @kropot72 @ganeshie8

OpenStudy (astrophysics):

Just wondering, is there more information, you have to make an equation.

ganeshie8 (ganeshie8):

if 2 represents the sample(100 bears) then 50 may represent the population( 100*25 bears )?

OpenStudy (astrophysics):

Will require cross multiplication, we can let a = bear population \[\frac{ a }{ 50 } = \frac{ 100 }{ 2 }\]

OpenStudy (rock_mit182):

there is no more info..

OpenStudy (rock_mit182):

so this is the property of extreme means ?

OpenStudy (rock_mit182):

\[2*a = 50*100\]

OpenStudy (rock_mit182):

\[a = \frac{ 50*100 }{ 2}\]

OpenStudy (rock_mit182):

\[a = 2500\]

OpenStudy (mathmate):

@rock_mit182 Just curious, are you working on hypergeometric distributions?

OpenStudy (mathmate):

@rock_mit182 are you there?

OpenStudy (mathmate):

@rock_mit182 This problem seems to be a typical hypergeometric distribution example with size of population 919 or 920.

OpenStudy (rock_mit182):

as a matter of fact.. are you a biologist ?

OpenStudy (mathmate):

I wish! lol Anyway, if you want to pursue the hypergeometric distribution route, let me know.